(3.) A fair six-sided die is rolled repeatedly. Let R denote the random variable representing the outcome of any particular roll. The following random variables are all discrete-time Markov chains. Specify the transition probabilities for each (as a check, make sure the row sums equal 1) (a) Xn, which represents the largest number obtained by the nth roll. (b) Yn, which represents the number of sixes obtained in n rolls.
Answer #1

Since R represents the random variable representing the outcome of any particular roll , hence R can take any value from 1 to 6

i.e R= {1,2,3,4,5,6}

(a) Let Xn-1 represents the largest number till (n-1)th roll and Xn represents the largest number till nth roll.

Then if Xn-1 = 1 , then Xn can take any value from 1 to 6 with equal probability 1/6 depending on the number obtained in the nth roll.

If Xn-1 = 2, then Xn cannot take 1 since 2>1. Hence Xn will be 2 if nth roll shows 1 or 2 but will be 3,4,5 or 6 if the dice shows 3,4,5 or 6 respectively. Hence Xn =2 with probability 2/6 and Xn = 3,4,5,6 with probability 1/6

If Xn-1 = 3, then Xn cannot take 1,2 since 3>1,2 . Hence Xn will be 3 if nth roll shows 1,2 or 3 but will be 4,5 or 6 if the dice shows 4,5 or 6 respectively. Hence Xn =3 with probability 3/6 and Xn = 4,5,6 with probability 1/6

Similarly proceeding if If Xn-1 = 6, then Xn cannot take any value other than 6 since 6>1,2,3,4,5. Hence Xn will be 6 with probability 1.

Hence our transition probability matrix in this case is

Xn

P = Xn-1  0 0 0 4/6 1/6 1/6 0 0 0 0 5/6 1/6

where (i,j)th element represents the probability of reaching state j in the nth step given the process is in state i in the (n-1)th step. For example 3/6 represents the probability that the process will be in state 3 in the nth role given it was in state 3 in (n-1) th roll . Similarly 2/6 represents the probability that the process will be in state 2 in the nth role given it was in state 2 in (n-1) th roll.  

(b) Let Yn-1 represents the number of sixes obtained till (n-1)th roll and Yn represents the number of sixes obtained till nth roll.

Then if Yn-1 = 0 (0 sixes obtained till (n-1)th roll), then Yn =1 with probability 1/6 (if 6 is obtained in nth roll) and Yn = 0 with probability 5/6 (if 6 is not obtained)

Similarly if Yn-1 = 1 (1 six obtained till (n-1)th roll), then Yn =2 with probability 1/6 (if 6 is obtained in nth roll) and Yn = 1 with probability 5/6 (if 6 is not obtained) .

Proceeding in this way we get  if Yn-1 = n-1 (0 sixes obtained till (n-1)th roll), then Yn =n with probability 1/6 (if 6 is obtained in nth roll) and Yn = n-1 with probability 5/6 (if 6 is not obtained) .

So our transition probability matrix is given by